### v2

# Towards Highly Realistic Artistic Style Transfer via Stable Diffusion with Step-aware and Layer-aware Prompt

Abstract
Artistic style transfer aims to transfer the learned artistic style onto an arbitrary content image, generating artistic stylized images. Existing generative adversarial network-based methods fail to generate highly realistic stylized images and always introduce obvious artifacts and disharmonious patterns. Recently, large-scale pre-trained diffusion models opened up a new way for generating highly realistic artistic stylized images. However, diffusion model-based methods generally fail to preserve the content structure of input content images well, introducing some undesired content structure and style patterns. To address the above problems, we propose a novel pre-trained diffusion-based artistic style transfer method, called LSAST, which can generate highly realistic artistic stylized images while preserving the content structure of input content images well, without bringing obvious artifacts and disharmonious style patterns. Specifically, we introduce a Step-aware and Layer-aware Prompt Space, a set of learnable prompts, which can learn the style information from the collection of artworks and dynamically adjusts the input images’ content structure and style pattern. To train our prompt space, we propose a novel inversion method, called Step-ware and Layer-aware Prompt Inversion, which allows the prompt space to learn the style information of the artworks collection. In addition, we inject a pre-trained conditional branch of ControlNet into our LSAST, which further improved our framework’s ability to maintain content structure. Extensive experiments demonstrate that our proposed method can generate more highly realistic artistic stylized images than the state-of-the-art artistic style transfer methods. Code is available at https://github.com/Jamie-Cheung/LSAST.
Both authors contributed equally to this research.    Corresponding authors.
1 Introduction
Artistic style transfer has recently attracted widespread attention in academia and industry since the seminal work of CycleGAN Zhu et al. (2017). Existing artistic style transfer methods can be divided into generative adversarial network-based approaches (GAN-based approaches) and large-scale pre-trained diffusion model-based approaches (Diffusion-based approaches).
Figure 1: Stylization examples with three different styles (i.e., Van Gogh, Monet, Gauguin). Compared to existing state-of-the-art large-scale pre-trained diffusion model-based methods (b-d) and generative adversarial network-based methods (e-h), our proposed method (a) generates highly realistic artistic stylized images and preserves the content structure of input content images well.
More specifically, GAN-based methods Zhu et al. (2017); Sanakoyeu et al. (2018); Kim et al. (2019); Park et al. (2020); Chen et al. (2023); Zhang et al. (2023d); Li et al. (2023c) generally utilize generative adversarial network and a training set of aligned/unaligned image pairs to learn the mapping between an input image and an output image. For example, Zhu et al. Zhu et al. (2017) used two mirror generative adversarial network to learn and improve the mapping between the input image and output image, synthesizing artistic stylized images. Sanakoyeu et al. Sanakoyeu et al. (2018) proposed a style-aware content loss, to improve the quality of artistic stylized images by capturing how style patterns affect content structure. However, GAN-based methods are limited by the instability of adversarial training and the scarcity of training data, failing to generate highly realistic artistic stylized images and introducing the obvious artifacts and disharmonious patterns on the stylized images (Please see in Fig. 1 (e-h)).
Large-scale pre-trained diffusion model-based approaches Nichol et al. (2021); Wu (2022); Ho et al. (2020); Hu et al. (2023); Li et al. (2024) use massive parameters to learn and store the information from the large-scale training data, possessing the ability to generate highly realistic images. This opens up a new way for generating highly realistic artistic stylized images. For example, Zhang et al. Zhang et al. (2024b) proposed a global prompt space, a learnable parameter matrix, to learn and store the style information from the collection of artworks and condition pre-trained large-scale diffusion model to generate artistic stylized image. Zhang et al. Zhang et al. (2023b) introduced a step-aware prompt space, a set of learnable parameter matrixes, for the whole diffusion process to generate desired stylized image. Although these approaches could generate highly realistic stylizied images, they failed to preserve the content structure of input content image well, bringing some undesired content structure and style patterns (Please see in Fig. 1 (b-d)).
To summarize, the main contribution of this paper is as follows:
  • We propose a novel pre-trained diffusion-based artistic style transfer framework, which can generate highly realistic stylized images and preserve the content structure of input content images well without introducing obvious artifacts and disharmonious style patterns.
  • We design a novel Step-aware and Layer-aware Prompt Space and conduct a Step-aware and Layer-aware Prompt Inversion to train prompt space, which can learn the style information from the collection of artworks and dynamically adjusts the input images’ content structure and style pattern.
  • Extensive quantitative and qualitative experiments demonstrate that our proposed LSAST outperforms the state-of-the-art GAN-based and Diffusion-based methods.
2 Related Work
Generative Adversarial Network-based Methods.
Generative adversarial network-based methods refer to the use of discriminator networks Zhu et al. (2017) to train a well-designed forward network that can bridge a mapping between the input image and output image. As the seminal work of artistic style transfer, Zhu et al. Zhu et al. (2017) expanded adversarial loss and proposed Cycle-Consistent loss to improve the quality of artistic stylized images. The cycle-consistent loss inspired a lot of researchers to explore a more effective way to enhance further the quality of artistic stylized images, including Wang et al. (2022); Park et al. (2020); Zheng et al. (2021); Fu et al. (2019); Zuo et al. (2023); Zhang et al. (2023d, 2021). For example, Park et al. Park et al. (2020) proposed a method to preserve the content structure of input content image by maximizing the mutual information between input image and output image via contrastive learning Zhang et al. (2024a). Sanakoyeu et al. Sanakoyeu et al. (2018) proposed a style-aware content loss to improve the stylization of images by capturing how style affects content. Zheng et al. Zheng et al. (2021) exploited the spatial patterns of self-similarity to capture spatial relationships within an image rather than domain appearance to preserve the content structure. While generative adversarial network-based methods effectively transfer the learned style onto an arbitrary content image, they fail to generate highly realistic artistic stylized images, introducing obvious artifacts and disharmonious patterns.
Acknowledgments
This work was supported in part by National Social Science Foundation Major Project “Research on Virtual Restoration of Tang and Song Painting Colors and Construction of Traditional Color Resource Library from a Digital Perspective” (19ZDA046), Zhejiang Province Program (2022C01222, 2023C03199, 2023C03201), the National Program of China (62172365, 2021YFF0900604, 19ZDA197), Ningbo Science and Technology Plan Project (2022Z167, 2023Z137), and MOE Frontier Science Center for Brain Science & Brain-Machine Integration (Zhejiang University).